How much work does the electric field do in moving a -8.0 μC charge from ground to a point whose potential is +70 V higher?
Answer #1

Here,

electric charge , Q = - 8 uC

work done by electric field = -Q * V

work done by electric field = 8 *10^-6 * 70

work done by electric field = 560 *10^-6 J

work done by electric field = 5.6 *10-4 J